OpenStudy (anonymous):

Help please :c Ralph wants to find the solution to a system of equations using y = x - 2 and y = x2 + 6x - 38. Joey says that Ralph can solve x2 + 5x - 36 = 0 to find the x coordinates of the solutions to his system. Explain and demonstrate why Joey is correct.

OpenStudy (anonymous):

@zepdrix any idea?

zepdrix (zepdrix):

So what Joey did was realize that the `solutions` to the system are where they equal one another,\[\Large\rm x-2=x^2+6x-38\]

zepdrix (zepdrix):

Understand how to proceed from there?

OpenStudy (anonymous):

Yes, you would add 2 correct?

zepdrix (zepdrix):

add 2, subtract x, to get the 0 on the left.

OpenStudy (anonymous):

x+6-36=0?

OpenStudy (anonymous):

x-30=0

zepdrix (zepdrix):

what? +_+

OpenStudy (anonymous):

x=30?

OpenStudy (anonymous):

I suck at math I'm sorry :c

zepdrix (zepdrix):

Subtracting x from each side, x-x removes the x on the left side,\[\Large\rm -2=x^2+6x-38-x\]Adding 2 to each side, -2+2 cancel out as well on the left,\[\Large\rm 0=x^2+6x-38-x+2\]Then combine like terms, 6x-x -38+2

OpenStudy (anonymous):

0+x^2+6x^2-36?

OpenStudy (anonymous):

0=***

zepdrix (zepdrix):

No you didn't combine the x terms.

OpenStudy (anonymous):

Would I add or subtract the x from 6x?

zepdrix (zepdrix):

subtract..

OpenStudy (anonymous):

Pshh I knew that.. 0=x^2+6-36

zepdrix (zepdrix):

No no no. You don't combine like terms like that. If x is an apple Then we have 6apples minus 1apple. Leaves us with 5apples ( or 5 x's ). 6x-x = 5x

OpenStudy (anonymous):

Oh.. Like I said I'm terrible at math..

OpenStudy (anonymous):

So 0=x^2+5x-36

zepdrix (zepdrix):

\[\Large\rm 0=x^2+5x-36\]Mmm ok good. And if we solve this, we can show that Joey's method in fact works.

OpenStudy (anonymous):

Okay, so we'd solve this by using the quadratic formula right?

zepdrix (zepdrix):

We could, but this one factors quite nicely.

zepdrix (zepdrix):

You want values that `multiply` to -36, but also `add` to positive 5.

zepdrix (zepdrix):

Let's try a few combinations. \[\Large\rm -6\cdot 6=36,\qquad\qquad -6+6\ne 5\]Hmm those factors didn't work.

zepdrix (zepdrix):

Hmm this combination is also no good. \[\Large\rm -18\cdot2=36,\qquad\qquad -18+2\ne5\]

zepdrix (zepdrix):

Woops those should say -36 in the examples*

OpenStudy (anonymous):

12 and 3?

zepdrix (zepdrix):

12-3 = 9 Hmm those don't give us 5 either. Can you think of any other factors of 39? :o Think of your 9's table.

zepdrix (zepdrix):

of 36*

OpenStudy (anonymous):

9 and 4?

OpenStudy (anonymous):

9*4=36

zepdrix (zepdrix):

Mmm ok let's try those. In order that we get a -36, one of them needs to be negative. You want the 9 or 4 to be negative.

OpenStudy (anonymous):

The 4 :oo

zepdrix (zepdrix):

Mmm\[\Large\rm -4\cdot9=-36,\qquad\qquad -4+9=5\qquad \color{green}{\checkmark}\]

zepdrix (zepdrix):

Ok great we found our factors!

OpenStudy (anonymous):

so you'd end up with 0=(x+9)(x-4)

OpenStudy (anonymous):

Oh okay. So x=9 and x=-4? :oo

OpenStudy (anonymous):

Or would it be switched around so 9 would be negative and 4 would be positive?

zepdrix (zepdrix):

I dunno why I said it would be subtraction. You had it right the first time -_- Sorry bout that, I must be tired... 0=(x+9)(x-4)

OpenStudy (anonymous):

okay.. So my answer would be x=-9 and x=4 correct? And don't worry about it, putting up with my stupidity would make me tired too xD

zepdrix (zepdrix):

Mmmm yah good job \c:/

OpenStudy (anonymous):

Thank you so much! c: